OpenStudy (anonymous):

Imagine that you are outside one sunny day and are standing 10 feet away from a tree that is 8 feet tall. Which function correctly represents the angle θ you make with the tree? θ=sin−1(8/10) θ=cos−1(8/10) θ=tan−1(8/10) θ=sec−1(8/10)

OpenStudy (mathteacher1729):

Have you tried sketching a picture?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):



OpenStudy (anonymous):

@welshfella

OpenStudy (mayankdevnani):

so can you write a relation between theta(angle) , Perpendicular and base of right angled triangle ?

OpenStudy (mayankdevnani):

HINT :- Use Trigonometry

OpenStudy (anonymous):

yea this is trigonometry

OpenStudy (mayankdevnani):

so write a relation between angles and sides ?

OpenStudy (anonymous):

now what is the relation perpendicular to the base

OpenStudy (anonymous):

sin? cos? tan? sec?

OpenStudy (mayankdevnani):

https://www.mathsisfun.com/sine-cosine-tangent.html

OpenStudy (mayankdevnani):

REMEMBER THIS :- http://www.mathwords.com/s/sohcahtoa.htm

OpenStudy (anonymous):

tan=opp/adj

OpenStudy (mayankdevnani):

correct !

OpenStudy (mayankdevnani):

so, \[\large \bf \tan \theta=\frac{8}{10}\]

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

now, take inverse of it

OpenStudy (mayankdevnani):

to get angle THETA

OpenStudy (mayankdevnani):

@DaWizjr

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

How do I take the inverse

OpenStudy (mayankdevnani):

\[\large \bf \tan \theta=\frac{8}{10}\] to take inverse, just shift tan to right hand side with its power negative,that's inverse so, \[\large \bf \theta=\tan^{-1} \frac{8}{10}=\color{red}{Answer}\]

OpenStudy (anonymous):

yea that makes since

OpenStudy (anonymous):

thanks

OpenStudy (mayankdevnani):

but this is not the actual method for taking inverse, i said incompletely because you are very small in the eyes of trig. So, please don't ask me about INVERSE,

OpenStudy (anonymous):

I might need you for more look out for me

OpenStudy (mayankdevnani):

sure , if i can

OpenStudy (anonymous):

Yea I know its always a another complicated way

OpenStudy (mayankdevnani):

yup !